The H+ concentration in an aqueous solution at 25°C is 4.3 x 10-4 What is [OH-]? 1 2 3 Х + +- 0 x 100 x 100
Answer #1

Given:
[H+] = 4.3*10^-4 M

use:
[OH-] = Kw/[H+]
Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC
[OH-] = (1.0*10^-14)/[H+]
[OH-] = (1.0*10^-14)/(4.3*10^-4)
[OH-] = 2.326*10^-11 M
Answer: 2.3*10^-11 M